the average age of four sisters is 24. The average age of eldest three is 26, how old is the youngest sister?
Found 3 solutions by ankor@dixie-net.com, greenestamps, ikleyn:
Answer by ankor@dixie-net.com(22740)

let y = the youngest sister age
:
The average age of four sisters is 24.
3s + y = 4(24)
3s + y = 96
The average age of eldest three is 26,
3s = 3(26)
3s = 78
how old is the youngest sister?
3s + y = 96
replace 3s with 78
78 + y = 96
y = 96 - 78
y = 18 yrs is the youngest sister

Answer by greenestamps(13200)



If an algebraic solution is not required, this can be solved informally using the fact that all the ages are "close together".

The average age of all four is 24; the average age of the three oldest is 26. On average, then, the age of each of the three oldest is 2 years above the average.

So all together, the ages of those three is 3*2=6 over the average. That means the age of the youngest has to be 6 years under the average.

ANSWER: 24-6 = 18

Answer by ikleyn(52781)

.
For such simple problem, than simpler the solution is, than better you understand it.


If the average age of four sisters is 24 years, then the sum of their ages is 4*24 = 96 years.


If the average age of the three oldest sisters is 26, then the sum of their ages is 3*26 = 78 years.


If the age of the youngest sister is x, then you have this simple equation


    x + 78 = 96,


from which you get the ANSWER  momentarily


    x = 96 - 78 = 18.

Solved.

When you learn on how to solve such problems, you will be able do it in one line, as THIS :

    ANSWER = 4*24 - 3*26 . . . 

--------------------

To get more practice in solving problems on average values, look into the lessons
    - Solved problems on average scores, weight, height and temperature
    - Solved problems on average scores
    - Solved problems on average age
in this site.

Consider these lessons as your textbook,  handbook,  tutorials and  (free of charge)  home teacher.